in a study about the hourly wage paid to young people aged 14-18 with a holiday job, the following summarised results were measured For a random sample of 400 young people. Ê400^Xi=3296 Ê400^Xi2^= 27833.35. a) calculate the mean and standard deviation of the data. b) use 95% confidence interval to answer the question. c) perform a hypothesis test at the 5% level of significance to answer the question. d) do you need to know why X is normally distributed?
1
Expert's answer
2020-03-26T16:46:14-0400

Given that

"\sum_{\mathclap{1\le i\le 400}} X_i=3296, \ \ \ \ \sum_{\mathclap{1\le i\le 400}} X_i^2=27833.35"

a) calculate the mean and standard deviation of the data


"\mu=E(X)={1 \over n}\sum_{\mathclap{1\le i\le n}} X_i={1 \over 400}\sum_{\mathclap{1\le i\le 400}} X_i="

"={3296 \over n}=8.24"


"Var(X)=\sigma^2=E(X^2)-(E(X))^2="

"={27833.35 \over 400}-(8.24)^2=1.685775"

"\sigma=\sqrt{\sigma^2}=\sqrt{1.685775}\approx1.2984"


b)We need to construct the "95\%" confidence interval for the population mean "\mu" with unknown population variance.

The critical value for "\alpha=0.05" and "df=n-1=399" degrees of freedom is


"t_c=t_{1-\alpha/2;n}=1.966"

The corresponding confidence interval is computed as shown below:


"CI=\big(\bar{X}-t_c\times{s \over \sqrt{n}},\ \bar{X}+t_c\times{s \over \sqrt{n}}\big)="

"=\big(8.24-1.966\times{1.2984 \over \sqrt{400}},\ 8.24+1.966\times{1.2984 \over \sqrt{400}}\big)="


"=(8.112,\ 8.368)"

If we use z-test.

The critical value for "\alpha=0.05" is "z_c=z_{1-\alpha/2}=1.96"

The corresponding confidence interval is computed as shown below:


"CI=\big(\bar{X}-z_c\times{\sigma \over \sqrt{n}},\ \bar{X}+z_c\times{\sigma \over \sqrt{n}}\big)="

"=\big(8.24-1.96\times{1.2984 \over \sqrt{400}},\ 8.24+1.96\times{1.2984 \over \sqrt{400}}\big)="

"=(8.113,\ 8.367)"

c) The following null and alternative hypotheses need to be tested:

"H_0: \mu=8"

"H_1: \mu\not=8"

This corresponds to a two-tailed test, for which a t-test for one mean, with unknown population standard deviation will be used.

Based on the information provided, the significance level is "\alpha=0.05," and the critical value for a two-tailed test is "t_c=1.966."  

The rejection region for this two-tailed test is "R=\{t:|t|>1.966\}."  

The t-statistic is computed as follows:


"t={\bar{X}-\mu_0 \over s/\sqrt{n}}={8.24-8 \over 1.2984/\sqrt{400}}\approx3.697"

Since it is observed that "|t|=3.697>1.966=t_c," it is then concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu" is different than "8," at the "0.05" significance level.


Using the P-value approach: The p-value for "t=3.697, \alpha=0.05" is "p= 0.000249." Since "p= 0.000249<0.05," it is concluded that the null hypothesis is rejected. Therefore, there is enough evidence to claim that the population mean "\mu" is different than "8," at the "0.05" significance level.


d) When the sample size is large, the "z-"tests are easily modified to yield valid test procedures without requiring either a normal population distribution or known standard deviation.